# Interpreting CLIP’s Image Representation via Text-Based Decomposition

### v1

@@ 3.1 CLIP-ViT Preliminaries
Given a batch of images and corresponding text descriptions $\{(I_{i},t_{i})\}_{i\in\{1,...,k\}}$, CLIP is trained to maximize the similarity of the image representation $M_{\text{image}}(I_{i})$ to its corresponding text representation $M_{\text{text}}(t_{i})$, while minimizing $\operatorname{sim}(M_{\text{image}}(I_{i}),M_{\text{text}}(t_{j}))$ for every $i\neq j$ in the batch.
@@ 3.3 Fine-grained decomposition into heads and positions
where $W_{VO}^{l,h}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ are transition matrices and $\alpha_{i}^{l,h}\in\mathbb{R}$ are the attention weights from the class token to the $i$-th token ($\sum_{i=0}^{N}{\alpha_{i}^{l,h}}=1$). Therefore, the MSA output can be decomposed into direct effects of individual heads and tokens.
@@
$\sum_{l=1}^{L}{P\left[\mathsf{MSA}^{l}(T^{l-1})\right]_{cls}}=\sum_{l=1}^{L}{\sum_{h=1}^{H}{\sum_{i=0}^{N}{{c_{i,l,h}}}}},\;~{}~{}c_{i,l,h}=Px_{i}^{l,h}$ (6)
@@ 4.1 Text-interpretable decomposition into heads
       $D\leftarrow RC^{T}$ $j^{*}\leftarrow\operatorname*{arg\,max}_{j=1}^{M}\mathrm{Var}(D[j])$ $\mathcal{T}\leftarrow\mathcal{T}\cup\{t_{j^{*}}\}$ for k in $[1,...,K]$ do
             $C^{\prime}[k]\leftarrow C^{\prime}[k]+\frac{\langle C[k],R[j^{*}]\rangle}{\mid\mid R[j^{*}]\mid\mid^{2}}R[j^{*}]$ $C[k]\leftarrow C[k]-\frac{\langle C[k],R[j^{*}]\rangle}{\mid\mid R[j^{*}]\mid\mid^{2}}R[j^{*}]$
@@ 4.2 Experiments
Some attention heads capture specific image properties. We report selected head descriptions from TextSpan ($m=60$) in Table 1, with full results in Appendix A.4. For some heads, the top descriptions center around a single image property like texture (L23H12), shape (L22H1), object count (L23H10), and color (L22H11). This suggests that these heads capture specific image properties. We qualitatively verify that the text tracks these image properties by retrieving the images with the largest similarity $\langle M_{\text{text}}(t_{i}),c_{\text{head}}^{l,h}\rangle$ for the top extracted text descriptions $t_{i}$. The results in Fig. 5 and 11 show that the returned images indeed match the text.
@@
Figure 6: Top-4 nearest neighbors per head and image. We retrieve the most similar images to an input image by computing the similarity of the direct contributions of individual heads. As some heads capture specific aspects of the image (e.g. colors/objects), retrieval according to this metric results in images that are most similar regarding these aspects. See additional results in the appendix (Fig. 12).
@@ 5 Decomposition into Image Tokens
Quantitative segmentation results. We follow a standard protocol for evaluating heatmap-based explainability methods (Chefer et al., 2021). We first compute image heatmaps given descriptions of the image class (e.g. “An image of a {class}”) [^3] To normalize out bias terms, we subtract from the heatmap an averaged heatmap computed across all class descriptions in ImageNet.. We then binarize them (by applying a threshold) to obtain a foreground/background segmentation. We compare the segmentation quality to zero-shot segmentations produced by other explainability methods in the same manner.
@@ A.3 Text Descriptions
Table 5: The ChatGPT prompt for general image descriptions generation.
A.4 TextSpan outputs for CLIP-ViT-L
@@
A.5 Qualitative results for image token decomposition
@@
A.6 Most similar images to TextSpan results

### v4

@@ 3.1 CLIP-ViT Preliminaries
Given a batch of images and corresponding text descriptions $\{(I_{i},t_{i})\}_{i\in\{1,...,k\}}$, CLIP is trained to maximize the similarity of the image representation $M_{\text{image}}(I_{i})$ to its corresponding text representation $M_{\text{text}}(t_{i})$, while minimizing $\operatorname{sim}(I_{i},t_{j})$ for every $i\neq j$ in the batch.
@@ 3.3 Fine-grained decomposition into heads and positions
where $W_{VO}^{l,h}\in\mathbb{R}^{d\times d}$ are transition matrices and $\alpha_{i}^{l,h}\in\mathbb{R}$ are the attention weights from the class token to the $i$-th token ($\sum_{i=0}^{N}{\alpha_{i}^{l,h}}=1$). Therefore, the MSA output can be decomposed into direct effects of individual heads and tokens.
@@
$\sum_{l=1}^{L}{P\left[\mathsf{MSA}^{l}(Z^{l-1})\right]_{cls}}=\sum_{l=1}^{L}{\sum_{h=1}^{H}{\sum_{i=0}^{N}{{c_{i,l,h}}}}},\;~{}~{}c_{i,l,h}=Px_{i}^{l,h}$ (6)
@@ 4.1 Text-interpretable decomposition into heads
       $D\leftarrow RC^{T}$
      $j^{*}\leftarrow\operatorname*{arg\,max}_{j=1}^{M}\mathrm{Var}(D[j])$
      $\mathcal{T}\leftarrow\mathcal{T}\cup\{t_{j^{*}}\}$
      for k in $[1,...,K]$ do
             $C^{\prime}[k]\leftarrow C^{\prime}[k]+\frac{\langle C[k],R[j^{*}]\rangle}{\mid\mid R[j^{*}]\mid\mid^{2}}R[j^{*}]$
            $C[k]\leftarrow C[k]-\frac{\langle C[k],R[j^{*}]\rangle}{\mid\mid R[j^{*}]\mid\mid^{2}}R[j^{*}]$
@@ 4.2 Experiments
Some attention heads capture specific image properties. We report selected head descriptions from TextSpan ($m=60$) in Table 1, with full results in Appendix A.5. For some heads, the top descriptions center around a single image property like texture (L23H12), shape (L22H1), object count (L23H10), and color (L22H11). This suggests that these heads capture specific image properties. We qualitatively verify that the text tracks these image properties by retrieving the images with the largest similarity $\langle M_{\text{text}}(t_{i}),c_{\text{head}}^{l,h}\rangle$ for the top extracted text descriptions $t_{i}$. The results in Fig. 5 and 12 show that the returned images indeed match the text.
@@
Figure 6: Top-4 nearest neighbors per head and image. We retrieve the most similar images to an input image by computing the similarity of the direct contributions of individual heads. As some heads capture specific aspects of the image (e.g. colors/objects), retrieval according to this metric results in images that are most similar regarding these aspects. See additional results in the project page and appendix (Fig. 13).
@@ 5 Decomposition into Image Tokens
Quantitative segmentation results. We follow a standard protocol for evaluating heatmap-based explainability methods (Chefer et al., 2021). We first compute image heatmaps given descriptions of the image class (e.g. “An image of a {class}”) [^3] We normalize out bias terms by subtracting from the heatmap an averaged heatmap computed across all class descriptions in ImageNet.. We then binarize them (by applying a threshold) to obtain a foreground/background segmentation. We compare the segmentation quality to zero-shot segmentations produced by other explainability methods in the same manner.
@@ A.3 Text Descriptions
Class-specific text descriptions. We generate additional class-specific text descriptions, by prompting ChatGPT with the prompt template provided in table 5. We queried to model for each of the ImageNet class names. This process resulted in 28767 unique sentences.
@@
Class-specific text descriptions prompt
Provide 40 image characteristics that are true for almost all the images of {class}. Be as general as possible and give short descriptions presenting one characteristic at a time that can describe almost all the possible images of this category. Don’t mention the category name itself (which is “{class}”). Here are some possible phrases: “Image with texture of …”, “Picture taken in the geographical location of…”, ”Photo that is taken outdoors”, “Caricature with text”, “Image with the artistic style of…”, “Image with one/two/three objects”, “Illustration with the color palette …”, “Photo taken from above/below”, “Photograph taken during … season”. Just give the short titles, don’t explain why, and don’t combine two different concepts (with “or” or “and”).
Table 5: ChatGPT prompts for image descriptions generation.
A.4 Additional Initial Description Pool Ablation
We present additional ablation of the initial set of text descriptions provided to TextSpan. The text description generation processes for each of the pools are described in Section A.3.
As shown in Figure 8, using the class-specific descriptions pool that includes around $\times$8 more examples than the general descriptions pool, allows us to obtain higher accuracy with fewer descriptions per head (smaller $m$). Nevertheless, using each of the two pools results in relatively similar accuracy with $m=60$.
Figure 8: ImageNet classification accuracy for the image representation projected to TextSpan bases (additional results). We evaluate our algorithm for different initial description pools, and with different output sizes.
A.5 TextSpan outputs for CLIP-ViT-L
@@
A.6 Qualitative results for image token decomposition
@@
A.7 Most similar images to TextSpan results